Express 700.1400.2100.2800.3500.42000 in terms of factorial


1
Expert's answer
2022-06-06T08:37:11-0400
"700\cdot1400\cdot2100\cdot2800\cdot3500\cdot42000"

"=700(1)\cdot700(2)\cdot700(3)\cdot700(4)\cdot700(5)\cdot700(6)"

"=(700)^6(1\cdot2\cdot3\cdot4\cdot5\cdot6)"

"=(700)^6\cdot6!"